# Covariance test and universal bootstrap by operator norm

### v1

@@ Abstract
Testing covariance matrices is crucial in high-dimensional statistics. Traditional methods based on the Frobenius and supremum norms are widely used but often treat the covariance matrix as a vector and neglect its inherent matrix structure. This paper introduces a new testing framework based on the operator norm, designed to capture the spectral properties of the covariance matrix more accurately. The commonly used empirical bootstrap and multiplier bootstrap methods are shown to fail for operator norm-based statistics. To derive the critical values of this type of statistics, we propose a universal bootstrap procedure, utilizing the concept of universality from random matrix theory. Our method demonstrates consistency across both high- and ultra-high-dimensional regimes, accommodating scenarios where the dimension-to-sample-size ratio $p/n$ converges to a nonzero constant or diverges to infinity. As a byproduct, we provide the first proof of the Tracy-Widom law for the largest eigenvalue of sample covariances with non-Gaussian entries as $p/n\to\infty$. We also show such universality does not hold for the Frobenius norm and supremum norm statistics. Extensive simulations and a real-world data study support our findings, highlighting the favorable finite sample performance of the proposed operator norm-based statistics.
62H15,
62F40,
60B20,
Covariance test,
high-dimensional hypothesis test,
operator norm,
random matrix theory,
universality,
bootstrap,
keywords:
[class=MSC]
keywords:
[…]
, and
@@ 1 Introduction
Testing large covariance has received considerable attention due to its critical role in high-dimensional statistics. Generally, statistics for testing one-sample high-dimensional covariance matrices are based on the distance between the sample covariance and the hypothesized covariance matrices. As highlighted by Chen, Qiu and Zhang (2023), this distance is typically measured using two types of norms: the Frobenius norm (e.g., Chen, Zhang and Zhong (2010); Cai and Ma (2013)) and the supremum norm (e.g., Jiang (2004); Cai and Jiang (2011)), similar to their application in high-dimensional mean tests Chen and Qin (2010); Cai, Liu and Xia (2014).
In contrast to the Frobenius norm and the supremum norm which treat the covariance matrices as vectors, the operator norm captures the spectral structure of the covariance matrices and has recently gained significant attention. For $n$ independent, not necessarily identically distributed random vectors $\bm{X}_{1},\cdots,\bm{X}_{n}\in\mathbb{R}^{p}$ with zero mean and a common covariance matrix $\bm{\Sigma}$, consider the statistics
@@
where $\bm{\hat{\Sigma}}=\sum_{i=1}^{n}\bm{X}_{i}\bm{X}_{i}^{T}/n$ is the sample covariance matrix, $\bm{\Sigma}_{0}$ is the hypothesized covariance matrix, and the operator norm of a matrix $\bm{A}$ is defined as $\|\bm{A}\|_{\text{op}}=\sup_{\|\bm{u}\|=1}\|\bm{A}\bm{u}\|$. The statistic $T$ is fundamental in principal component analysis, as it bounds the error of sample eigenvalues and eigenvectors. As a result, numerous studies have developed non-asymptotic upper bounds for the tail of $T$, e.g., Adamczak et al. (2011); Bunea and Xiao (2015); Koltchinskii and Lounici (2017). However, these upper bounds are often conservative when used to construct confidence intervals for $T$. Only a few studies have explored the asymptotic distribution of $T$ due to the complex nature of the spectral analysis, including Han, Xu and Zhou (2018); Lopes (2022a); Lopes, Erichson and Mahoney (2023); Giessing (2023). Nonetheless, these works assume that either the dimension $p$ grows slower than $n$, or that the eigenvalues of $\bm{\Sigma}$ and $\bm{\Sigma}_{0}$ decay fast enough, making its effective rank smaller than $n$. Under such conditions, the empirical covariance matrix $\bm{\hat{\Sigma}}$ is a consistent estimator of $\bm{\Sigma}$ with respect to the operator norm. Essentially, these assumptions reduce the intrinsic dimension of the problem, leaving the true high-dimensional cases unresolved. This article studies the behavior of $T$ under both high- and ultra-high-dimensional regimes, characterized by constants $C_{1},C_{2}>0$ and $\alpha\geq 1$, such that
$C_{1}n^{\alpha}\leq p\leq C_{2}n^{\alpha}.$ (2)
Furthermore, no eigen-decay assumptions are made for $\bm{\Sigma}$ or $\bm{\Sigma}_{0}$. In this context, the dimension-to-sample size ratio $\phi=p/n$ may converge to a nonzero constant or diverge to infinity, and no consistent estimators exist for $\bm{\Sigma}$ under the operator norm, as discussed in Ding and Wang (2023); Ding, Hu and Wang (2024). This setting, while challenging, is common in practice. For example, the simple case of $\bm{\Sigma}=\bm{I}_{p}$ with $p$ growing proportionally to $n$ falls into this inconsistent regime. Consequently, existing results are not applicable to such cases. The theoretical difficulties in determining the distribution of $T$ restrict the use of the operator norm in high-dimensional covariance testing within the inconsistent regime.
1.1 Random matrix theory
Building upon prior work, various approaches have been developed for analyzing $T$ and its variants in specific scenarios using random matrix theory. When $\bm{\Sigma}_{0}=\bm{I}_{p}$, the distribution of $T$ can be derived by examining the limiting behavior of the extreme eigenvalues of the sample covariance matrix $\bm{\hat{\Sigma}}$. The study of extreme spectral properties of high-dimensional covariance matrices has attracted substantial interest, resulting in a number of influential contributions, including Johnstone (2001); Bianchi et al. (2011); Onatski, Moreira and Hallin (2013); Johnstone and Paul (2018). Building on the progress of these studies, El Karoui (2007); Bao, Pan and Zhou (2015); Lee and Schnelli (2016); Knowles and Yin (2017) established that under conditions where the dimension $p$ and sample size $n$ grow proportionally and other certain regularity conditions, the limiting distribution of the largest eigenvalue of $\bm{\hat{\Sigma}}$ follows the Tracy–Widom law. In cases where $\bm{\Sigma}_{0}$ is invertible, Bao, Pan and Zhou (2015) proposed transforming the data $\bm{X}_{1},\cdots,\bm{X}_{n}$ into $\bm{\Sigma}_{0}^{-1/2}\bm{X}_{1},\cdots,\bm{\Sigma}_{0}^{-1/2}\bm{X}_{n}$, allowing for testing whether the covariance of $\bm{\Sigma}_{0}^{-1/2}\bm{X}_{i}$ is the identity matrix $\bm{I}_{p}$. This transformation results in the statistic $T$ taking the form of Roy’s largest root statistic
$T^{\text{Roy}}=\left\|\bm{\Sigma}_{0}^{-\frac{1}{2}}\bm{\hat{\Sigma}}\bm{\Sigma}_{0}^{-\frac{1}{2}}-\bm{I}_{p}\right\|_{\text{op}},$ (3)
and the Tracy–Widom distribution results hold for $T^{\text{Roy}}$ when $p$ and $n$ grow proportionally, i.e., $\alpha=1$. However, similar distributional results for $T$ have not yet been obtained for general $\bm{\Sigma}_{0}$, even when $p/n$ is bounded. To address this gap, we leverage random matrix theory to characterize the extreme singular values for the matrix of the form $\hat{\bm{\Sigma}}+\bm{R}$ with a general matrix $\bm{R}$. When taking $\bm{R}=-\bm{\Sigma}_{0}$, this gives the limiting properties of $T$. Additionally, we introduce a novel technique, the universal bootstrap, designed to directly approximate the distribution of $T$ in the context of general $\bm{\Sigma}_{0}$.
1.2 Bootstrap
@@
The bootstrap method, a generic resampling technique to approximate the distribution of a statistic, was first introduced by Efron (1979). Originally designed for fixed-dimensional problems, the bootstrap method has recently been adapted for high-dimensional settings through substantial work. Chernozhukov, Chetverikov and Kato (2013, 2017); Lopes, Lin and Müller (2020); Lopes (2022b); Chernozhukov, Chetverikov and Koike (2023) developed Gaussian approximation methods and established the approximation rates for empirical bootstrap and multiplier bootstrap for the supremum norm of the sum of high-dimensional vectors. Moreover, Han, Xu and Zhou (2018); Lopes, Blandino and Aue (2019); Yao and Lopes (2021); Lopes (2022a); Lopes, Erichson and Mahoney (2023) explored bootstrap methods for spectral statistics of covariance matrices, while these works often assume a low intrinsic dimension or effective rank relative to the sample size $n$. Despite such advances in high-dimensional bootstrap methods, El Karoui and Purdom (2019) and Yu, Zhao and Zhou (2024) demonstrated that empirical and multiplier bootstraps can be inconsistent for $T$, even when $\bm{\Sigma}_{0}=\bm{I}_{p}$ if $p$ and $n$ grow proportionally and the eigenvalues of $\bm{\Sigma}$ and $\bm{\Sigma}_{0}$ do not decay.
To address the inconsistency issues associated with high-dimensional bootstraps based on Gaussian approximation, motivated by the concept of universality, a topic of recent interest in random matrix theory (Hu and Lu, 2022; Montanari and Saeed, 2022), we develop a novel resampling method informed by universality principles. Specifically, traditional bootstrap methods (Chernozhukov, Chetverikov and Kato, 2013, 2017; Chernozhukov, Chetverikov and Koike, 2023), depending on the high-dimensional central limit theorem, treat the large sample covariance matrix $\bm{\hat{\Sigma}}=\sum_{i=1}^{n}\bm{X}_{i}\bm{X}_{i}^{T}/n$ as a sum of random matrices and approximate its distribution with a Gaussian matrix sharing the same covariance as $\bm{\hat{\Sigma}}$. Nonetheless, as the dimension $p$ increases, the accuracy of this approximation diminishes, eventually leading to inconsistency, as shown by El Karoui and Purdom (2019). This inconsistency indicates that in high-dimensional contexts, the large matrix structures dominate the central limit theorem’s applicability, causing the sample covariance matrix $\bm{\hat{\Sigma}}$ to diverge from Gaussian behavior. To circumvent this limitation, the universal property leverages high-dimensional structures. Broadly, the universality suggests that the asymptotic distribution of $T$ only depends on the distribution of $\bm{X}_{1},\cdots,\bm{X}_{n}$ through their first two moments. This allows us to construct the universal bootstrap statistic by substituting independent Gaussian samples $\bm{Y}_{1},\dots,\bm{Y}_{n}\sim\mathcal{N}(\bm{0},\bm{\Sigma})$ in place of $\bm{X}_{1},\dots,\bm{X}_{n}$ in the definition of (1),
@@
where $\bm{\hat{\Sigma}}^{\text{ub}}=\sum_{i=1}^{n}\bm{Y}_{i}\bm{Y}_{i}^{T}/n$. The key insight here is that while the universal bootstrap matrix $\bm{\hat{\Sigma}}^{\text{ub}}$ need not approximate a Gaussian matrix, it effectively replicates the structure of $\bm{\hat{\Sigma}}$, which is disrupted by the empirical and multiplier bootstraps. Although this method relies on sophisticated random matrix theory, the implementation remains straightforward.
@@ 1.3 Our contributions
We summarize our contribution as three-fold. First, we establish the anisotropic local law for matrices of the form $\hat{\bm{\Sigma}}+\bm{R}$ within both high- and ultra-high-dimensional settings (2), where $\bm{R}$ may have positive, negative, or zero eigenvalues. Erdős, Yau and Yin (2012); Knowles and Yin (2017) demonstrated that the anisotropic local law is essential for proving eigenvalue rigidity and universality. The anisotropic local law for $\hat{\bm{\Sigma}}$ has been previously established by Knowles and Yin (2017) for $\alpha=1$ and by Ding and Wang (2023); Ding, Hu and Wang (2024) for $\alpha>1$. However, the existence of $\bm{R}$ alters the structure of the sample covariance matrix $\hat{\bm{\Sigma}}$, rendering existing techniques invalid. Additionally, due to the potential ultra-high dimension, each block of the Green function matrix (see (21) for definition) may converge at different rates, making the bound derived in Ding and Wang (2023) suboptimal. To address these challenges, we introduce a new ”double Schur inversion” technique to represent the Green function matrix using local blocks. We also define an auxiliary parameter-dependent covariance matrix and establish the corresponding parameter-dependent Marčenko-Pastur law. This parameter-dependent structure enables us to prove the entry-wise local law, which represents a weaker version of the anisotropic local law. Furthermore, we present an improved representation of the anisotropic local law, enhancing the results of Ding and Wang (2023) and Ding, Hu and Wang (2024). This unified structure provides a valuable tool for establishing the anisotropic local law.
Second, we establish the universality results for the statistic $T$, based on which the universal bootstrap procedure is introduced. Leveraging this universality, we demonstrate that the empirical distribution generated by the universal bootstrap in (4) effectively approximates the distribution of $T$. Additionally, the asymptotic distribution of $T$ is shown to be independent of the third and fourth moments of $\bm{X}_{1},\cdots,\bm{X}_{n}$, allowing us to focus primarily on the covariance. In contrast, we find that two commonly used norms, the Frobenius norm and supremum norm, lack this universality, as their asymptotic distributions explicitly depend on all fourth moments of the data. Therefore, standardized statistics based on these norms requires fourth-moment estimation, which is generally more complex than estimating the covariance. This difference highlights a key advantage of employing the operator norm for covariance testing.
@@
Third, we perform size and power analyses for the statistics $T$ and $T^{\text{Roy}}$ and propose a combined approach to enhance testing power. We also develop a generalized universality theorem to show the consistency of the universal bootstrap for statistics based on extreme eigenvalues, including this combined statistic. For size analysis, we demonstrate that the universality results apply to $T^{\text{Roy}}$. As a byproduct, we extend the Tracy–Widom law for the largest eigenvalue of sample covariance with general entries under the ultra-high dimension regime, addressing a long-standing gap since it was only proven for Gaussian entries by Karoui (2003). For power performance, we analyze both $T$ and $T^{\text{Roy}}$ within the generalized spiked model framework from Bai and Yao (2012); Jiang and Bai (2021). We show that $T^{\text{Roy}}$ performs better in worst-case scenarios while $T$ excels on average performance. To further enhance the power, we propose a new combined statistic, $T^{\text{Com}}$, supported by a generalized universality theorem that confirms the validity of the universal bootstrap for $T^{\text{Com}}$. This also serves as a theoretical guarantee for the universal bootstrap applicable to a broader range of other statistics based on extreme eigenvalues. Extensive simulations validate these findings, highlighting the superior performance of our combined statistics across a range of scenarios.
@@ 1.4 Notations and paper organization
Throughout the paper, we reserve boldfaced symbols for vectors and matrices. For a complex number $z\in\mathbb{C}$, we use $\Re z$ and $\Im z$ for its real part and imaginary part, respectively. For a vector $\bm{u}\in\mathbb{R}^{p}$, we use $\|\bm{u}\|=\sqrt{\sum_{i=1}^{p}u_{i}^{2}}$ for its Euclidean norm. For a matrix $\bm{A}=(a_{ij})_{M\times N}$, we use $\|\bm{A}\|_{\text{op}}$, $\|\bm{A}\|_{\text{F}}=:\sqrt{\sum_{i=1}^{M}\sum_{j=1}^{N}a_{ij}^{2}}$, and $\|\bm{A}\|_{\text{sup}}=:\sup_{1\leq i\leq M,1\leq j\leq N}|a_{ij}|$ to represent its operator norm, Frobenius norm and the supremum norm, respectively. Denote the singular values of $\bm{A}$ by $\sigma_{1}(\bm{A})\geq\sigma_{2}(\bm{A})\geq\cdots\geq\sigma_{r}(\bm{A})$, where $r=\min\{M,N\}$. When $M=N$, we use $\lambda_{1}(\bm{A}),\lambda_{2}(\bm{A}),\cdots\lambda_{M}(\bm{A})$ for the eigenvalues of $\bm{A}$. The trace of a square matrix $\bm{A}$ is denoted by $\text{tr}(\bm{A})=\sum_{i=1}^{p}a_{ii}$. For two sequences $\{a_{n}\}_{n=1}^{\infty}$, $\{b_{n}\}_{n=1}^{\infty}$, we use $a_{n}\lesssim b_{n}$ or $a_{n}=O(b_{n})$ to show there exists a constant $C$ not depending on $n$ such that $|a_{n}|\leq Cb_{n}$ for all $n\in\mathbb{N}$. Denote $a_{n}=o(b_{n})$ if $a_{n}/b_{n}\to 0$, and $a_{n}\asymp b_{n}$ if both $a_{n}\lesssim b_{n}$ and $b_{n}\lesssim a_{n}$.
@@
The paper is organized as follows. Section 2 provides an overview of the proposed method and presents our main results on the consistency of the proposed universal bootstrap procedure. Section 3 describes the key tools from random matrix theory and applies the universality theorem to covariance testing problems. Section 4 presents the power analysis of the universal bootstrap procedure. Section 5 provides simulation results and a real data example, demonstrating the numerical performance of our operator norm-based statistics. Detailed technical proofs are provided in the supplementary material. The data and codes are publicly available in a GitHub repository (https://github.com/zhang-guoyu/universal_bootstrap).
2 Proposed method and theoretical outlines
In this section, we introduce the covariance testing problem and present an informal summary of our main universal results. The formal results are provided in Section 3.3. We also discuss statistical applications of our proposed universal bootstrap method, which utilizes the concept of universality. Consider $n$ independent random vectors $\bm{X}_{1},\cdots,\bm{X}_{n}\in\mathbb{R}^{p}$ with zero mean and covariance matrix $\bm{\Sigma}=(\sigma_{ij})_{p\times p}$, which are not necessarily identically distributed. For a given $p$ by $p$ non-negative matrix $\bm{\Sigma}_{0}$, we aim to test the hypothesis
$H_{0}:\bm{\Sigma}=\bm{\Sigma}_{0}\quad\text{vs.}\quad H_{1}:\bm{\Sigma}\neq\bm{\Sigma}_{0}.$ (5)
As we focus on testing the covariance structure, we allow the third and fourth moments of $\bm{X}_{i}$ to differ across $i=1,\cdots,n$.
To simplify notation, we arrange the data into an $n\times p$ matrix $\bm{X}=(\bm{X}_{1},\cdots,\bm{X}_{n})^{T}\in\mathbb{R}^{n\times p}$, where $\bm{A}^{T}$ denotes the transpose of $\bm{A}$, and the sample covariance matrix is expressed by $\bm{\hat{\Sigma}}=\bm{X}^{T}\bm{X}/n$. Recall the definition of statistic $T$ in (1), we aim to control the size of our procedure by charaterizing the asymptotic distribution of $T$ under the null hypothesis $H_{0}$. Consider the Gaussian data matrix $\bm{Y}=(\bm{Y}_{1},\cdots,\bm{Y}_{n})^{T}\in\mathbb{R}^{n\times p}$, with $\bm{Y}_{1},\cdots,\bm{Y}_{n}\sim\mathcal{N}(\bm{0},\bm{\Sigma})$. Its sample covariance is defined accordingly as $\bm{\hat{\Sigma}}^{\text{ub}}=\bm{Y}^{T}\bm{Y}/n$. Throughout this article, we work under both the proportionally growing high-dimensional regime and the ultra-high dimensional regime (2). Under these regimes, the key quantity of interest to bound is as follows
where $\mathbf{B}_{\text{op}}(\bm{\Sigma}_{0},t)=\big{\{}\bm{N}\ :\|\bm{N}-\bm{\Sigma}_{0}\|_{\text{op}}\leq t\big{\}}$ represents operator norm ball with center $\bm{\Sigma}_{0}$ and the radius $t$.
A simplified form is given to illustrate our main results.
Result 1 (Informal).
We have
$\rho_{n}(\bm{\Sigma})\leq Cn^{-\delta}\to 0\quad\text{as}\quad n\to\infty,$ (7)
for some constants $C$ and $\delta>0$, where $\bm{\Sigma}$ is the covariance matrix of $\bm{X}$.
This result enables us to approximate the asymptotic distribution of $T$ using $T^{\text{ub}}$, as defined in (4). Under $H_{0}$, $\bm{\Sigma}=\bm{\Sigma}_{0}$ and the distribution of $T^{\text{ub}}$ does not contain any unknown quantities. This inspires the procedure of universal bootstrap. Given the analytical complexity of $T^{\text{ub}}$’s distribution, we generate $B$ independent samples, $\bm{Y}^{1},\dots,\bm{Y}^{B}$, from the same distribution as $\bm{Y}$. We define $\bm{\hat{\Sigma}}^{\text{ub},b}=(\bm{Y}^{b})^{T}(\bm{Y}^{b})/n$ for each $b=1,\dots,B$, allowing us to compute

### v2

# Universal Bootstrap for Spectral Statistics: Beyond Gaussian Approximation

@@ Abstract
Spectral analysis plays a crucial role in high-dimensional statistics, where determining the asymptotic distribution of various spectral statistics remains a challenging task. Due to the difficulties of deriving the analytic form, recent advances have explored data-driven bootstrap methods for this purpose. However, widely used Gaussian approximation-based bootstrap methods, such as the empirical bootstrap and multiplier bootstrap, have been shown to be inconsistent in approximating the distributions of spectral statistics in high-dimensional settings. To address this issue, we propose a universal bootstrap procedure based on the concept of universality from random matrix theory. Our method consistently approximates a broad class of spectral statistics across both high- and ultra-high-dimensional regimes, accommodating scenarios where the dimension-to-sample-size ratio $p/n$ converges to a nonzero constant or diverges to infinity without requiring structural assumptions on the population covariance matrix, such as eigenvalue decay or low effective rank. We showcase this universal bootstrap method for high-dimensional covariance inference. Extensive simulations and a real-world data study support our findings, highlighting the favorable finite sample performance of the proposed universal bootstrap procedure.
Keywords: covariance inference, operator norm, random matrix theory, universality
@@ 1 Introduction
Spectral analysis of large covariance matrices is crucial in multivariate statistical estimation and hypothesis testing. Consider $n$ independent, not necessarily identically distributed random vectors $\bm{X}_{1},\cdots,\bm{X}_{n}\in\mathbb{R}^{p}$ with zero mean and a common covariance matrix $\bm{\Sigma}$. A prominent example we consider is the spectral statistic
@@
where $\bm{\hat{\Sigma}}=\sum_{i=1}^{n}\bm{X}_{i}\bm{X}_{i}^{T}/n$ is the sample covariance matrix, $\bm{\Sigma}_{0}$ is the pre-specified hypothesized covariance matrix, and the operator norm of a matrix $\bm{A}$ is defined as $\|\bm{A}\|_{\text{op}}=\sup_{\|\bm{u}\|=1}\|\bm{A}\bm{u}\|$. The statistic $T$ is fundamental in high-dimensional statistics, including covariance estimation (Fan et al., 2011), principal component analysis (Lopes et al., 2023), and common factor determination (Yu, Zhao & Zhou, 2024). Consequently, many studies have established non-asymptotic upper bounds on the tail behavior of $T$ (e.g., Adamczak et al., 2011; Bunea & Xiao, 2015; Koltchinskii & Lounici, 2017). However, these bounds tend to be conservative when constructing confidence intervals for $T$. Few studies have examined the asymptotic distribution and bootstrap approximation of $T$ for general $\bm{\Sigma}$, $\bm{\Sigma}_{0}$ (Han et al., 2018; Lopes, 2022b; Lopes et al., 2023; Giessing, 2023). Nonetheless, these works assume that either the dimension $p$ grows at a slower rate than $n$, or that the eigenvalues of $\bm{\Sigma}$ and $\bm{\Sigma}_{0}$ decay fast enough, making its effective rank smaller than $n$. These assumptions enable Gaussian approximation but effectively lower the problem’s intrinsic dimension, leaving true high-dimensional cases unaddressed.
This article studies the behavior of $T$ and other general spectral statistics under both high- and ultra-high-dimensional regimes, allowing the dimension-to-sample size ratio $\phi=p/n$ to converge to a nonzero constant or diverge to infinity. Unlike previous studies, we impose no eigen-decay assumptions on $\bm{\Sigma}$ or $\bm{\Sigma}_{0}$. Under these conditions, standard high-dimensional Gaussian approximations become invalid. To address this gap, we establish a new universality principle for spectral statistics, which leads to a theoretically justified and computationally efficient bootstrap method for approximating their distributions.
1.1 A universality approach: overcoming Gaussianity
@@
The bootstrap method is a generic resampling technique to approximate the distribution of a statistic. Originally designed for fixed-dimensional problems, the bootstrap method has recently been adapted for high-dimensional settings through substantial work. Chernozhukov et al. (2013, 2017); Deng & Zhang (2020); Lopes et al. (2020); Lopes (2022a); Chernozhukov et al. (2023) established Gaussian approximation rates for empirical and multiplier bootstrap for the supremum norm of high-dimensional mean vectors. Moreover, Han et al. (2018); Lopes et al. (2019); Yao & Lopes (2021); Lopes (2022b); Lopes et al. (2023) explored bootstrap methods for spectral statistics of covariance matrices, typically assuming a low intrinsic dimension or effective rank relative to the sample size $n$.
Despite these advances, El Karoui & Purdom (2019); Dette & Rohde (2024) and Yu, Zhao & Zhou (2024) demonstrated that the Gaussian approximation-based bootstrap for spectral statistics fails when the intrinsic dimension is comparable to the sample size. These bootstrap methods approximate the distribution of the sample covariance matrix $\bm{\hat{\Sigma}}=\sum_{i=1}^{n}\bm{X}_{i}\bm{X}_{i}^{T}/n$ using a Gaussian matrix with the same covariance. Nonetheless, as the dimension $p$ grows, this approximation becomes less accurate and eventually inconsistent, as shown by El Karoui & Purdom (2019). This suggests that high-dimensional covariance matrix structures constrain the applicability of the central limit theorem.
To overcome this limitation, we leverage the universality property of high-dimensional covariance structures. Universality suggests that the asymptotic distribution of $T$ only depends on the distribution of $\bm{X}_{1},\cdots,\bm{X}_{n}$ through their first two moments. This allows us to construct the universal bootstrap statistic by substituting independent Gaussian samples $\bm{Y}_{1},\dots,\bm{Y}_{n}\sim\mathcal{N}(\bm{0},\bm{\Sigma})$ in place of $\bm{X}_{1},\dots,\bm{X}_{n}$ in the definition of (1),
@@
where $\bm{\hat{\Sigma}}^{\text{ub}}=\sum_{i=1}^{n}\bm{Y}_{i}\bm{Y}_{i}^{T}/n$. The key insight is that although the bootstrapped matrix $\bm{\hat{\Sigma}}^{\text{ub}}$ does not approximate a Gaussian matrix, it effectively preserves the structure of $\bm{\hat{\Sigma}}$, which empirical and multiplier bootstraps fail to maintain. Despite its reliance on intricate random matrix theory, this method remains computationally straightforward to implement. Further extensions for general spectral statistics beyond $T$ are also developed.
1.2 A motivating example: covariance inference
Our universal bootstrap method applies directly to spectral statistics in high-dimensional covariance tests. Testing large covariance has received considerable attention due to its critical role in high-dimensional statistics. Typically, hypothesis tests for high-dimensional covariance matrices $H_{0}:\bm{\Sigma}=\bm{\Sigma}_{0}$ rely on the norm between $\bm{\hat{\Sigma}}$ and $\bm{\Sigma}_{0}$, commonly the Frobenius norm (e.g., Chen et al., 2010; Cai & Ma, 2013) or the supremum norm (e.g., Jiang, 2004; Cai & Jiang, 2011), paralleling their high-dimensional mean tests Chen & Qin (2010); Cai et al. (2014). In contrast to the Frobenius norm and the supremum norm, which treat the covariance matrices as vectors and neglect their spectral structures, the operator norm captures the spectral structure and has recently gained significant attention. Various approaches have been developed for analyzing $T$ and its variants in specific scenarios for covariance tests. When $\bm{\Sigma}_{0}=\bm{I}_{p}$, the distribution of $T$ can be derived by the extreme eigenvalues of the sample covariance matrix $\bm{\hat{\Sigma}}$. Karoui (2007); Bao et al. (2015); Lee & Schnelli (2016); Knowles & Yin (2017) established the limiting Tracy–Widom law for top eigenvalue of $\bm{\hat{\Sigma}}$ when dimension $p$ and sample size $n$ grow proportionally. In cases where $\bm{\Sigma}_{0}$ is invertible, Bao et al. (2015) proposed transforming the data $\bm{X}_{1},\cdots,\bm{X}_{n}$ into $\bm{\Sigma}_{0}^{-1/2}\bm{X}_{1},\cdots,\bm{\Sigma}_{0}^{-1/2}\bm{X}_{n}$ and considered the Roy’s largest root statistic
$T^{\text{Roy}}=\left\|\bm{\Sigma}_{0}^{-\frac{1}{2}}\bm{\hat{\Sigma}}\bm{\Sigma}_{0}^{-\frac{1}{2}}-\bm{I}_{p}\right\|_{\text{op}}.$ (3)
The limiting Tracy–Widom distribution also holds for $T^{\text{Roy}}$. However, for general $\bm{\Sigma}_{0}$, distributional results for $T$ remain unknown, even when $p/n$ is bounded. The universal bootstrap method addresses this gap by leveraging universality in random matrix theory. More applications in high-dimensional statistics are presented in Section 4.
@@ 1.3 Our contributions
We summarize our contribution as three-fold. First, we propose a universal bootstrap procedure for $T$ based on universality results. We show that this procedure effectively approximates distributions of spectral statistics. As a byproduct, we establish the Tracy–Widom law for the largest eigenvalue of sample covariance matrices with general entries in the ultra-high-dimensional regime, addressing a long-standing gap since Karoui (2003). Additionally, the asymptotic distribution of $T$ is shown to be independent of the third and fourth moments of entries. In contrast, the Frobenius norm and supremum norm lack this universality. Consequently, statistics based on Frobenius and supremum norms require fourth-moment estimation, which is generally more complex and computationally intensive. This underscores a key advantage of using operator norm-based spectral statistics.
@@
Second, we develop a unified bootstrap method for general spectral statistics and apply it to the covariance inference. We conduct power analysis for $T$ and $T^{\text{Roy}}$ within the generalized spiked model framework from Jiang & Bai (2021). We show that $T^{\text{Roy}}$ performs better in worst-case scenarios while $T$ excels in average performance. To further enhance the power, we propose a new combined statistic, $T^{\text{Com}}$, supported by a generalized universality result that serves as a theoretical guarantee for the universal bootstrap applicable to general spectral statistics. Extensive simulations validate these findings, highlighting the superior performance of our combined statistics across a wide range of scenarios.
Third, from a technical perspective, we establish the anisotropic local law and universality for matrices of the form $\hat{\bm{\Sigma}}+\bm{R}$ in both high- and ultra-high-dimensional settings. When $\bm{R}=-\bm{\Sigma}_{0}$, this result characterizes the limiting properties of $T$. The anisotropic local law for $\bm{R}=\bm{0}$ was previously established by Knowles & Yin (2017) for bounded $p/n$ and by Ding & Wang (2023); Ding et al. (2024) for diverging $p/n$. However, the presence of $\bm{R}$ changes the problem structure, and the ultra-high-dimensional setting makes existing bounds suboptimal. We address this gap by introducing a new weighted anisotropic local law, refining the results of Ding & Wang (2023) and Ding et al. (2024). This advancement strengthens the framework for establishing universality.
@@ 1.4 Notations and paper organization
Throughout the paper, we reserve boldfaced symbols for vectors and matrices. For a complex number $z\in\mathbb{C}$, we use $\Re z$ and $\Im z$ for its real part and imaginary part, respectively. For a vector $\bm{u}\in\mathbb{R}^{p}$, we use $\|\bm{u}\|=\sqrt{\sum_{i=1}^{p}u_{i}^{2}}$ for its Euclidean norm. Denote the singular values of a matrix $\bm{A}$ by $\sigma_{1}(\bm{A})\geq\sigma_{2}(\bm{A})\geq\cdots\geq\sigma_{r}(\bm{A})$, where $r=\text{rank}(\bm{A})$. We use $\lambda_{1}(\bm{A}),\lambda_{2}(\bm{A}),\cdots\lambda_{M}(\bm{A})$ for the eigenvalues of a squared $\bm{A}$. For two sequences $\{a_{n}\}_{n=1}^{\infty}$, $\{b_{n}\}_{n=1}^{\infty}$, we use $a_{n}\lesssim b_{n}$ or $a_{n}=O(b_{n})$ to show there exists a constant $C$ not depending on $n$ such that $|a_{n}|\leq Cb_{n}$ for all $n\in\mathbb{N}$. Denote $a_{n}=o(b_{n})$ if $a_{n}/b_{n}\to 0$.
@@
The paper is organized as follows. Section 2 provides an overview of the proposed method. Section 3 presents the main theoretical results on the consistency of the universal bootstrap procedure. Section 4 discusses the applications of our universal bootstrap in the covariance test and construction of simultaneous confidence intervals. Section 5 provides simulation results and a real data example, demonstrating the numerical performance of our methods. Detailed technical proofs, as well as the codes and datasets for reproducibility, are provided in the supplementary material.
2 Universal bootstrap
In this section, we present the universal bootstrap procedure and its generalized version. Consider $n$ independent random vectors $\bm{X}_{1},\cdots,\bm{X}_{n}\in\mathbb{R}^{p}$ with zero mean and covariance matrix $\bm{\Sigma}=(\sigma_{ij})_{p\times p}$, which are not necessarily identically distributed. As we focus on the covariance structure, we allow the third and fourth moments of $\bm{X}_{i}$ to differ across $i=1,\cdots,n$. To simplify notation, we arrange the data into an $n\times p$ matrix $\bm{X}=(\bm{X}_{1},\cdots,\bm{X}_{n})^{T}\in\mathbb{R}^{n\times p}$, where $\bm{A}^{T}$ denotes the transpose of $\bm{A}$, and the sample covariance matrix is expressed by $\bm{\hat{\Sigma}}=\bm{X}^{T}\bm{X}/n$.
Recalling the definition of statistic $T$ in (1), we aim to characterize the asymptotic distribution of $T$. Consider the Gaussian data matrix $\bm{Y}=(\bm{Y}_{1},\cdots,\bm{Y}_{n})^{T}\in\mathbb{R}^{n\times p}$, with $\bm{Y}_{1},\cdots,\bm{Y}_{n}\sim\mathcal{N}(\bm{0},\bm{\Sigma})$. Its sample covariance is defined accordingly as $\bm{\hat{\Sigma}}^{\text{ub}}=\bm{Y}^{T}\bm{Y}/n$. In Section 3, we demonstrate that the asymptotic distribution of $T$ can be uniformly approximated by $T^{\text{ub}}$, as defined in (2). This motivates the universal bootstrap procedure. In practice, due to the analytical complexity of the distribution of $T^{\text{ub}}$, we generate $B$ independent samples, $\bm{Y}^{1},\dots,\bm{Y}^{B}$, drawn from the same distribution as $\bm{Y}$. We compute the sample covariance matrix $\bm{\hat{\Sigma}}^{\text{ub},b}=(\bm{Y}^{b})^{T}(\bm{Y}^{b})/n$ for each $b=1,\dots,B$ and define